OpenStudy (anonymous):

let W be the line in R^2 with equation y=2x find equation for orthogonal complement of W

OpenStudy (anonymous):

I think it'd just be \(y=\frac{-1}{2}x \)

OpenStudy (anonymous):

how sir please explain

OpenStudy (anonymous):

you need the vectors that are orthogonal to the original vector

OpenStudy (anonymous):

which vectors are orthogonal to each other in this case please show i am not getting point how is -1/2x is orthogonal to 2x

OpenStudy (anonymous):

also how to find same thing for W is a plane in R^3 with equation x-2y-3z=0

OpenStudy (anonymous):

for lines to be perpendicular the product of the gradients must be -1 ( ie they are negative reciprocals )

OpenStudy (anonymous):

also how to find same thing for W is a plane in R^3 with equation x-2y-3z=0